OpenStudy (anonymous):

Given the parent function f(x) = 2x, which graph shows f(x) + 1?

OpenStudy (anonymous):

exponential function going through point 0, 2 and ending up on the right exponential function going through point 0, 0 and ending down on the right exponential function going through point 0, 5 and ending up on the right exponential function going through point 0, 1.5 and ending up on the right

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

first off i assume thisis \(f(x)=2^x\) since it says "exponential"

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then it looks just like \(u=y=2^x\) only shifted up one unit if \(x=0\) then \(f(0)=2^0+1=1+1=2\) so it has the point \((0,2)\) on the graph

OpenStudy (anonymous):

So it would be A?

OpenStudy (anonymous):

yes